The density of mercury is 13.6 g/mL. If a sample of mercury weighs 267g, what is

the

volume of the sample in milliliters?


1
Expert's answer
2021-10-28T02:57:52-0400

"ρ= \frac{m}{V} \\
13.6 = \frac{267}{V} \\
V = \frac{267}{13.6} = 19.63 \;mL"